# Distributionally Robust System Level Synthesis With Output Feedback Affine Control Policy

### v1

@@ 4.2 Characterization of Distribution Shift
To properly formulating and solving the DR-SLS problem in (33), the distribution shift between $\mathbb{P}^{\mathcal{M}}$ and $\overline{\mathbb{P}}^{\hat{\mathcal{M}}}$, which is measured by the Wasserstein distance $d_{W}(\overline{\mathbb{P}}^{\hat{\mathcal{M}}},\mathbb{P}^{\mathcal{M}})$ and the ambiguity set $\mathcal{B}_{\varepsilon}(\overline{\mathbb{P}}^{\hat{\mathcal{M}}})$, should be properly considered. An excessively large radius $\varepsilon$ of the ambiguity set will lead to a conservative solution, and an excessively small one would result in an aggressive and brittle solution. Besides, the distribution shift is influenced by the closed-loop model mismatch between $\mathcal{M}$ and $\hat{\mathcal{M}}$, which indicates its correlation with the closed-loop parameterization $\{\hat{\Phi}_{y},\hat{\Phi}_{u},\hat{\phi}_{y},\hat{\phi}_{u}\}$. Consequently, simply bounding the Wasserstein distance with a constant does not reflect this property and can lead to unsatisfactory performance.
@@
where the above relationships hold owning to that the innovation process $\mathbf{e}$ is coupled with $\bm{\eta}$ only through $\Phi$ in (24) and (25), and the sub-multiplicativity property of 1-norm. This completes the proof. $\square$
@@
$|\displaystyle\quad+||\hat{\Phi}R_{\Phi}(\tilde{\mathbf{y}}_{0}+\tilde{\Theta}\mathbf{e}_{i})||\bigg{)}+||(R_{\phi}-I)\hat{\phi}$$|$
@@ 4.3 Tractable Reformulation
While the DR-SLS formulation in (33) could enhance the robustness of the optimal solution of the SLS design, it is challenging to solve the corresponding optimization problem when considering the Wasserstein distance (41). In this section, a tractable reformulation is proposed. To this end, the following assumptions are assumed.
Assumption 2: There exists constants $\alpha>1$ and $\beta>0$ such that
@@
Term 1: $\hat{\Phi}(R_{\Phi}-I)(\hat{\mathbf{y}}_{0}+\hat{\Theta}\mathbf{e}_{i})||$. Based on the norm sub-multiplicativity, it can be derived
@@
Term 2: $\hat{\Phi}R_{\Phi}(\tilde{\mathbf{y}}_{0}+\tilde{\Theta}\mathbf{e}_{i})||$. Following a similar fashion in analyzing Term 1, and exploiting Assumption 4, it can be derived
@@ 5 Simulation Results
In this section, numerical simulation results are presented to demonstrate the effectiveness of the proposed approach. The numerical study is based on the linear system considered in [26], with a more challenging modification by setting $B_{2}=0$ such that small perturbations of $B_{2}$ will change the sign of the control input, resulting in different behaviours. The true system dynamics used in simulation is
@@
where $w_{t}\in\mathbb{R}^{2}$ and $v_{t}\in\mathbb{R}$ are assumed follow uniform distributions with each element belongs to $\mathcal{U}(-0.01,0.01)$. For constructing the innovation process, the feedback matrix $L$ in (1) is set as $0.1,0.1]^{\mathrm{T}}$, and ALCA-LC is Schur stable. For the SLS design, its prediction horizon is set as $T=15$, and $\tau=25$. The cost function is set as
@@
Fig. 2 presents the simulation results adopting the proposed DR-SLS approach. It can be observed from Fig. 2(a-b) that the output and input predictions all satisfy system constraints and leave margins to the respective permissible boundaries. In comparison with the case of the N-SLS approach shown in Fig. 1, the input-output predictions do not exactly converge to the origins, and small offsets are introduced to enhance robustness against possible model uncertainties. Fig. 2(c-d) show the true closed-loop trajectories. In contrast to the unstable behavior observed with the N-SLS approach, the DR-SLS approach ensures closed-loop stability while satisfying all system constraints, demonstrating its robustness and effectiveness.
@@
Figure 1: Output and input trajectories with N-SLS approach: (a) open-loop output prediction, (b) closed-loop output, (c) open-loop input prediction, (d) closed-loop input.
Figure 2: Output and input trajectories with DR-SLS approach: (a) open-loop output prediction, (b) closed-loop output, (c) open-loop input prediction, (d) closed-loop input.

### v2

@@ 4.2 Characterization of Distribution Shift
In order to properly formulate and solve the DR-SLS problem in (33), the distribution shift between $\mathbb{P}^{\mathcal{M}}$ and $\overline{\mathbb{P}}^{\hat{\mathcal{M}}}$, which is measured by the Wasserstein distance $d_{W}(\overline{\mathbb{P}}^{\hat{\mathcal{M}}},\mathbb{P}^{\mathcal{M}})$ and the associated ambiguity set $\mathcal{B}_{\varepsilon}(\overline{\mathbb{P}}^{\hat{\mathcal{M}}})$, should be properly considered. An excessively large radius $\varepsilon$ of the ambiguity set will lead to a conservative solution, and an excessively small one would result in an aggressive and fragile solution. Importantly, the Wasserstein distance $d_{W}(\overline{\mathbb{P}}^{\hat{\mathcal{M}}},\mathbb{P}^{\mathcal{M}})$ is determined by the closed-loop model mismatch between $\mathcal{M}$ and $\hat{\mathcal{M}}$, and is therefore a function of the predictive closed-loop parameterization $\{\hat{\Phi}_{y},\hat{\Phi}_{u},\hat{\phi}_{y},\hat{\phi}_{u}\}$. Consequently, simply setting the radius of the ambiguity set with a prescribed constant does not reflect its intrinsic dependence on the SLS parameters and can lead to unsatisfactory performance.
@@
where the above relationships are based on the fact that the innovation process $\mathbf{e}$ is coupled with $\bm{\eta}$ only through $\Phi$ in (24) and (25), and the sub-multiplicativity property of 1-norm. This completes the proof. $\square$
@@
$|\displaystyle\qquad+||\hat{\Phi}R_{\Phi}(\tilde{\mathbf{y}}_{0}+\tilde{\Theta}\mathbf{e}_{i})||\bigg)+||(R_{\phi}-I)\hat{\phi}$$|$
@@ 4.3 Tractable Reformulation
While the DR-SLS formulation in (33) could enhance the robustness of the optimal solution of the SLS design, it is challenging to solve the corresponding optimization problem when considering the Wasserstein distance (41). In this section, a tractable reformulation is proposed. To this end, the following assumptions are made.
Assumption 2: There exist constants $\alpha>1$ and $\beta>0$ such that
@@
Term 1: $\hat{\Phi}(R_{\Phi}-I)(\hat{\mathbf{y}}_{0}+\hat{\Theta}\mathbf{e}_{i})||$. The norm sub-multiplicativity can be exploited to show that
@@
Term 2: $\hat{\Phi}R_{\Phi}(\tilde{\mathbf{y}}_{0}+\tilde{\Theta}\mathbf{e}_{i})||$. Following a similar analysis as in Term 1, and exploiting Assumption 4, it can be derived that
@@ 5 Simulation Results
In this section, numerical simulation results are presented to demonstrate the effectiveness of the proposed approach. The numerical study is based on the linear system considered in [26], with a more challenging modification by setting $B_{2}=0$ such that small perturbations of $B_{2}$ will change the sign of the control input, resulting in different behaviours. The true system dynamics used in simulation are
@@
where $w_{t}\in\mathbb{R}^{2}$ and $v_{t}\in\mathbb{R}$ are assumed to follow uniform distributions with each element belonging to $\mathcal{U}(-0.01,0.01)$. For constructing the innovation process, the feedback matrix $L$ in (1) is set as $0.1,0.1]^{\mathrm{T}}$, and ALCA-LC is Schur stable. For the SLS design, its prediction horizon is set as $T=15$, and $\tau=25$. The cost function is set as
@@
Figure 1: Output and input trajectories with N-SLS approach: (a) open-loop output prediction, (b) open-loop input prediction, (c) closed-loop output, (d) closed-loop input.
Figure 2: Output and input trajectories with DR-SLS approach: (a) open-loop output prediction, (b) open-loop input prediction, (c) closed-loop output, (d) closed-loop input.
Fig. 2 presents the simulation results adopting the proposed DR-SLS approach. It can be observed from Fig. 2(a-b) that the output and input predictions all satisfy system constraints and leave margins to the respective permissible boundaries. In comparison with the case of the N-SLS approach shown in Fig. 1, the input-output predictions do not exactly converge to the origin, and small offsets are introduced to enhance robustness against possible model uncertainties. Fig. 2(c-d) show the true closed-loop trajectories. In contrast to the unstable behavior observed with the N-SLS approach, the DR-SLS approach ensures closed-loop stability while satisfying all system constraints, demonstrating its robustness and effectiveness.